OpenStudy (anonymous):

1/4 is to 1 and 1/4 as is 2 is to b.... solve for "b"

OpenStudy (anonymous):

Are you sure this is physics? This seems like it either should go under math or english because of the analogy...

OpenStudy (anonymous):

so you are asking, c*1/4=1+1/4 c*2=? solve for c in the first equation. plug it in the second to get an answer.

OpenStudy (stormfire1):

If I understand the question correctly...it's a simple ratio problem:\[\frac{\frac{1}{4}}{1\frac{1}{4}}=\frac{2}{b}\]\[\frac{1}{5}=\frac{2}{b}\]\[b=10\]